OpenStudy (anonymous):

The world's population is expected to grow at a rate of 1.3% per year until at least the year 2020. In 1994 the total population of the world was about 5,642,000,000 people. Use the formula Pn = P0ein to predict the world's population, Pn, n years after 1994, with P0 equal to the population in 1994 and i equal to the expected growth rate. What is the world's predicted population n the year 2020, rounded to the nearest million?

OpenStudy (anonymous):

Given: P₀ = 5,642×10^6 people, i = 1.3% = 0.013 per year, n = 2020 - 1994 = 26 years, We get: Pn = P₀ e^(i.n) . . .= 5,642×10^6×e^(0.013×26) . . .= 7,911×10^6 people

OpenStudy (anonymous):

Thanks!